inequality: 8 - x _> 6x + 29
expression: Simplify the expression
3(x-2)-9(x^2+7x +4)-5(x+8)

Answer by tracyc0912(40)

8-x>= 6x+29 (subtract 29 from each side)
-21-x>= 6x (add 1x to each side)
-21>=7x (divide each side by 7)
-3>=x
3(x-2) - 9(x^2+7x+4)-5(x+8) (distribute first)
3x-6-9x^2-63x-36-5x-40 (combine like terms)
-9x^2-65x-82
Good Luck :-)